# Randomized Physics-Informed Machine Learning for Uncertainty Quantification in High-Dimensional Inverse Problems

### 11 Dec 2023 v1

@@ 2.4.2 Metropolis rejection for the non-linear model
Next, we define a random vector $\bm{\delta}=\mathcal{R}(\mathbf{x};\bm{\xi}^{*},\bm{\eta}^{*})-\bm{\omega}$ and assume that there exists an invertible map $\mathcal{G}:(\bm{\omega},\bm{\alpha},\bm{\beta})\rightarrow(\bm{\delta},\bm{\xi}^{*},\bm{\eta}^{*})$, where $\bm{\xi}^{*}$ and $\bm{\eta}^{*}$ minimize the rPICKLE loss function $L^{r}$. Since $p(\bm{\omega},\bm{\alpha},\bm{\beta})$ is known, the joint distribution of $(\bm{\delta},\bm{\xi}^{*},\bm{\eta}^{*})$ can be computed as
@@
The first sample is automatically accepted. The acceptance or rejection of $\bm{\zeta}_{k}$ ($k>1$) sample can be done according to the independent Metropolis-Hastings (IMH) method Tierney (1994) with the acceptance ratio $\alpha$:
@@
It should be noted that in this IMH algorithm, the proposal function $q$ is not conditioned on the previous sample as in the HMC method. However, the acceptance ratio depends on the previous sample and the total transition obeys the Markov property. In the case of the linear $\mathcal{R}$ model considered in Section 2.4.1, every proposal is accepted because the determinant of the Jacobian is a constant and independent of $\bm{\zeta}$.
@@ 3.1 Governing equations
We use a cell-centered finite volume (FV) discretization for Eqs. (102) - (104) and the two-point flux approximation (TPFA) to compute residuals in the rPICKLE objective function. The numerical domain $\Omega$ is discretized into $N$ finite volume cells, and $u(\mathbf{x})$ and $y(\mathbf{x})$ are approximated with $\mathbf{u}=\{u_{1},...,u_{N}\}$ and $\mathbf{y}=\{y_{1},...,y_{N}\}$ vectors of their respective values evaluated at the cell centers $\{\mathbf{x}_{1},...,\mathbf{x}_{N}\}$. For the inverse problem, we assume there are $N^{\text{obs}}_{y}$ observations of $y$, $\mathbf{y}^{\text{obs}}=\{y_{i}^{\text{obs}}\}_{i=1}^{N_{y}^{\text{obs}}}$, and $N_{u}^{\text{obs}}$ observations of $u$, $\mathbf{u}^{\text{obs}}=\{u^{\text{obs}}_{i}\}_{i=1}^{N_{u}^{\text{obs}}}$.
@@ 3.2 Hanford Site Case Study
The rPICKLE method is tested for estimating subsurface parameters at the Hanford Site, a United States Department of Energy site situated on the Columbia River in Washington State. We use Eqs. (102)-(104) to describe the two-dimensional (depth-averaged) groundwater flow at the Hanford Site. The ground truth conductivity and transmissivity fields are taken from a previous calibration study reported in Cole et al. (2001). Following Yeung et al. (2022), we employ the unstructured quadrilateral mesh with $N=1475$ cells (see Figure 1) to compute PDE residuals. The rPICKLE estimates of the posterior distribution of log-transmissivity are compared with that of the HMC and, for the mode of the distribution, with PICKLE.
@@
To enable a comparison between HMC and rPICKLE, we generate a lower-dimensional (smoother) $y$ field $y_{\mathrm{LD}}$ using iterative local averaging of $y_{\mathrm{HD}}$. Such averaging reduces the variance and increases the correlation length of the field Tartakovsky et al. (2017) and, therefore, reduces the dimensionality of the inverse problem. At $k$th iteration, for FV element $i$, we calculate the geometric mean of $y^{(k)}_{j}$ over the adjacent $j$ elements and assign this value to $y^{(k+1)}_{i}$. Here, $y_{\mathrm{HD}}=y^{(k=0)}$ and $y_{\mathrm{LD}}=y^{(k=30)}$. We find that the ten-term CKLE of $y_{\mathrm{LD}}$ retains 95% of the total variance of this field.
@@
Figure 2: (a) The reference log-transmissivity field $y_{\mathrm{HD}}$ and (b) the reference hydraulic head field $u_{\mathrm{HD}}$ corresponding to $y_{\mathrm{HD}}$ obtained by solving Eqn (102) - (104) with the finite volume method. (c) The low-dimensional field $y_{\mathrm{LD}}$ obtained by iterative local averaging of $y_{\mathrm{HD}}$ and (d) the hydraulic head field $u_{\mathrm{LD}}$ corresponding to $y_{\mathrm{LD}}$.
@@ 3.3 Prior Mean and Covariance Models of $y$ and $u$
Next, we generate $N^{\text{MC}}$ number of realizations of the stochastic $y$ field by independently sampling $\{\xi_{i}\}_{i=1}^{N_{\xi}}$ from the normal distribution and solving Eqs. (102)-(104) for each realization of $y$ using the FV method described in Yeung et al. (2022). The ensemble of $u$ solutions is used to compute the (ensemble) mean and covariance of $u$ using Eqs. (112) and (113). Then, the mean and covariance of $u$ are conditioned on $u$ observations using the GPR equations (114). Finally, the CKLE of $u$ in Eq. (4) is constructed by performing the eigenvalue decomposition of the conditional covariance of $u$, i.e., by solving the eigenvalue problem similar to Eq. (111). In this work, we set $N^{\text{MC}}=5,000$.
@@ 4.1 Low-dimensional problem
We first present results for the low-dimensional case with the ground truth $y$ field given by $y_{\mathrm{LD}}$. Here, we assume that ten $y_{\mathrm{LD}}$ and $u_{\mathrm{LD}}$ observations are available. The locations of observations are randomly selected from the Hanford Site well locations. We use the HMC-sampled posterior distribution of $\bm{\xi}$ (and $y$), and the PICKLE-estimated MAP to benchmark the rPICKLE method with and without Metropolization. As stated earlier, we set $\sigma^{2}_{\xi}=\sigma^{2}_{\eta}=1$. The value of $\sigma^{2}_{r}$ is chosen to maximize the LPP of the posterior. However, the joint problem of minimizing the rPICKLE loss over $\bm{\xi}$ and $\bm{\eta}$ and maximizing the LPP over $\sigma^{2}_{r}$ is computationally challenging. Instead, we compute the inverse solutions for several values of $\sigma^{2}_{r}$ in the range $[10^{-5},1]$ and select the value producing the largest LPP. We also study the effect of $\sigma^{2}_{r}$ on uncertainty in the inverse solution and the performance of HMC and rPICKLE methods.
@@
In Figure 3, we also show the coordinates of the joint posterior mode given by the PICKLE solution. We can see that the coordinates of the modes of the marginal and bivariate distributions obtained from HMC and rPICKLE are in good agreement with the coordinates of the joint distribution mode. It should be noted that unless the posterior is Gaussian, the coordinates of the modes of marginal distributions and the corresponding coordinates of the joint distribution mode may not coincide. The fact that for the low-dimensional case, the coordinates of the marginal and joint distributions’ modes are close to each other indicates that the posterior distribution is close to Gaussian.
@@
Finally, we present the Bayesian predictions of $y$. For each of the three methods and $\sigma^{2}_{r}=10^{-2}$, Figure 4 shows the estimated mean $\mu_{\hat{y}}(\mathbf{x}|\mathcal{D}_{res})$ field, the absolute difference between the mean and the reference field $|y-\mu_{\hat{y}}(\mathbf{x}|\mathcal{D}_{res})|$, the standard deviation $\sigma_{\hat{y}}(\mathbf{x}|\mathcal{D}_{res})$, and the coverage plot showing the locations where the reference solution is within the $95\%$ confidence interval. In Table 1, we summarize results in terms of the relative $\ell_{2}$ error and $\ell_{\infty}$ error between the predicted mean and the reference $y_{\mathrm{LD}}$ fields, LPP, and the coverage (percentage of nodes where the reference solution is within the $95\%$ confidence interval) for $\sigma^{2}_{r}=10^{-5}$, $10^{-4}$, $10^{-2}$, $10^{-1}$, and $10^{0}$. For a given value of $\sigma^{2}_{r}$, all three methods produce a posterior mean close to the PICKLE’s estimate of MAP, and similar LPPs and coverages. This indicates that (i) the HMC and rPICKLE sampled distributions converge to the same posterior (note that in Section 2.4.1, we only prove the consistency of rPICKLE in the linear case), and (ii) Metropolization makes the posterior more descriptive of the reference field, i.e., it yields larger LPP and smaller $\ell_{2}$ error, but the improvements are less than $1\%$. We also find that, for this problem, the total acceptance rate for Metropolization is above $95\%$, which agrees with the findings in Oliver et al. (1996); Wang et al. (2018) that the acceptance rate in randomized algorithms is, in general, very high. It should be noted that the Metropolization requires the estimation of Jacobian, which becomes computationally expensive for high-dimensional problems. For this reason, in the high-dimensional case presented in Section 4.2, we do not perform Metropolization and accept all samples generated by the rPICKLE algorithm.
@@
According to Table 1, for a given value of $\sigma^{2}_{r}$, HMC and rPICKLE yield similar mean estimates of $y$ in terms of the point errors and the relative $\ell_{2}$ errors. Both, the HMC and rPICKLE errors are smallest for $\sigma^{2}_{r}=10^{-2}$. We find that the LPPs in these methods are also the largest for this value of $\sigma^{2}_{r}$. This indicates that $\sigma^{2}_{r}=10^{-2}$ provides the most informative posterior distribution of $y$. It is worth reminding that $\sigma^{2}_{r}=\gamma$, the regularization parameter in PICKLE. Since PICKLE’s $\ell_{2}$ and $\ell_{\infty}$ errors are also smallest for $\sigma^{2}_{r}=10^{-2}$, we conclude that this value of $\sigma^{2}_{r}$ also provides the optimal regularization for this problem.
@@
Figure 4: The low-dimensional $y$ field estimated from (a) HMC (b) rPICKLE and (c) rPICKLE with Metropolization. The first column – the posterior mean estimates; the second column – point errors in the predicted mean with respect to the reference $y_{LD}$ field; the third column – the posterior standard deviation of $y$; and the fourth column – the coverage of the reference field with the $95\%$ credibility interval, where zero and one values correspond to the reference field being outside and inside the credibility interval, respectively.
Table 1: The summary of the rPICKLE, HMC, and PICKLE results for the low-dimensional problem with different values of $\sigma^{2}_{r}$. The relative $\ell_{2}$ and $\ell_{\infty}$ errors in the estimated $y$ field with respect to the reference field, the log predictive probability (LPP), and the percentage of coverage of $y_{\mathrm{LD}}$ by the $95\%$ confidence interval. Note that the PICKLE solution only provides MAP, and LPP and coverage for the PICKLE solution are not provided.
@@ 4.2 High-dimensional problem
Here, we consider the IUQ problem where the $y$ field is given by $y_{\mathrm{HD}}$. First, we estimate the posterior distributions for $\sigma^{2}_{r}=10^{-4}$, $10^{-2}$, and $10^{-1}$ given $N_{y}^{\text{obs}}=100$ observations of the $y_{\mathrm{HD}}$ field. Later, we estimate posteriors for $N_{y}^{\text{obs}}=50$ and $200$ to study the dependence of posterior on the number of $y$ measurements. In all simulations in this section, we assume that $N_{u}^{\text{obs}}=323$, i.e., $u$ measurements are available at all wells.
@@
Table 2: The summary of the rPICKLE and PICKLE results for the high-dimensional problem for the priors based on different numbers of $y$ observations and $\sigma^{2}_{r}$. The relative $\ell_{2}$ and $\ell_{\infty}$ errors in the estimated $y$ field with respect to the reference field, the log predictive probability (LPP), and the percentage of coverage of $y_{\mathrm{LD}}$ by the $95\%$ confidence interval.
@@
Table 2 summarizes the relative $\ell_{2}$ and $\ell_{\infty}$ errors, LPP, and coverage in rPICKLE estimates of $y$ for $\sigma^{2}_{r}=10^{-4}$, $10^{-2}$, and $10^{-1}$. We find that the smallest rPICKLE errors and the largest LPP are achieved for $\sigma^{2}_{r}=10^{-2}$. However, we find that LPP is more sensitive to $\sigma^{2}_{r}$ than $\ell_{2}$ error – $\ell_{2}$ errors vary by less than 7% for the considered $\sigma^{2}_{r}$ values while LPP values change by more than 100%. Also, the $\ell_{2}$ error is 1% smaller for $\sigma^{2}_{r}=10^{-4}$ than for $\sigma^{2}_{r}=10^{-1}$ but the LPP is 70% larger for $\sigma^{2}_{r}=10^{-1}$ than for $\sigma^{2}_{r}=10^{-4}$. Therefore, we conclude that LPP provides a better criterion for selecting $\sigma^{2}_{r}$ than $\ell 2$ errors. Errors in the PICKLE and rPIKCLE predictions of $y$ are very similar, and $\sigma^{2}_{r}=10^{-2}$ also provides the best value of regularization coefficient for PICKLE, i.e., the PICKLE error is smallest for $\gamma=\sigma^{2}_{r}=10^{-2}$.
@@
Figure 6 shows the rPICKLE estimate of the posterior mean of $y$, the absolute point difference between the mean and reference $y_{\mathrm{HD}}(x)$ fields, the posterior standard deviation of $y(x)$, and the coverage for $\sigma^{2}_{r}=10^{-2}$ and $\sigma^{2}_{r}=10^{-4}$. We see significant differences in rPICKLE predictions for different $\sigma^{2}_{r}$. Errors in the predictions with $\sigma^{2}_{r}=10^{-4}$ are in general larger than in the predictions with $\sigma^{2}_{r}=10^{-2}$ with the maximum point error being 50% larger. As expected, the posterior standard deviations are in general larger in the prediction with the larger $\sigma^{2}_{r}$, however, the maximum point standard deviation is larger in the simulation with the smaller $\sigma^{2}_{r}$. We also see that $\sigma^{2}_{r}=10^{-2}$ produces a better coverage – the reference $y_{\mathrm{HD}}$ field is within the confidence interval in 82% of all predicted locations versus 65% for $\sigma^{2}_{r}=10^{-2}$. We remind that the LPP for $\sigma^{2}_{r}=10^{-4}$ is -2792, which is significantly smaller than that for $\sigma^{2}_{r}=10^{-2}$ (2032).
Next, we study uncertainty in inverse rPICKLE solution as a function of $N_{y}^{\text{obs}}$. Figure 7 depicts the estimates of $y$ obtained with rPICKLE for $N_{y}^{\text{obs}}=50$ and $200$ and $\sigma^{2}_{r}=10^{-4}$. The estimates for $N_{y}^{\text{obs}}=100$ are given in Figure 6. Table 2 summarizes the relative $\ell_{2}$ and $\ell_{\infty}$ errors, LPP, and the percent of coverage of the corresponding posteriors. As $N_{y}^{\text{obs}}$ increases, the posterior mean becomes closer to $y_{\mathrm{HD}}$, and the posterior variance of $y$ decreases. The LPP increases with $N_{y}^{\text{obs}}$ indicating that the posterior distribution becomes more informative. Also, we see that for all values of $N_{y}^{\text{obs}}$, the coverage for $y_{\mathrm{HD}}$ field is good, with the best coverage ($75\%$) achieved for $N_{y}^{\text{obs}}=200$.

### 23 Dec 2023 v2

@@ 2.4.2 Metropolis rejection
Next, we define a random vector $\bm{\delta}=\mathcal{R}(\mathbf{x};\bm{\xi}^{*},\bm{\eta}^{*})-\bm{\omega}$ and assume that there exists an invertible map $\mathcal{G}:(\bm{\omega},\bm{\alpha},\bm{\beta})\rightarrow(\bm{\delta},\bm{\xi}^{*},\bm{\eta}^{*})$, where $\bm{\xi}^{*}$ and $\bm{\eta}^{*}$ minimize the rPICKLE loss function $L^{r}$. Because $p(\bm{\omega},\bm{\alpha},\bm{\beta})$ is known, the joint distribution of $(\bm{\delta},\bm{\xi}^{*},\bm{\eta}^{*})$ can be computed as
@@
The first sample is automatically accepted. The acceptance or rejection of the $\bm{\zeta}_{k}$ ($k>1$) sample can be done according to the independent Metropolis-Hastings (IMH) method Tierney (1994) with the acceptance ratio $\alpha$:
@@
It should be noted that in this IMH algorithm, the proposal function $q$ is not conditioned on the previous sample as in the HMC method. However, the acceptance ratio depends on the previous sample, and the total transition obeys the Markov property. In the case of the linear $\mathcal{R}$ model considered in Section 2.4.1, every proposal is accepted because the determinant of the Jacobian is a constant and independent of $\bm{\zeta}$.
@@ 3.1 Governing Equations
We use a cell-centered finite volume (FV) discretization for Eqs. (102)–(104) and the two-point flux approximation (TPFA) to compute residuals in the rPICKLE objective function. The numerical domain $\Omega$ is discretized into $N$ finite volume cells, and $u(\mathbf{x})$ and $y(\mathbf{x})$ are approximated with $\mathbf{u}=\{u_{1},...,u_{N}\}$ and $\mathbf{y}=\{y_{1},...,y_{N}\}$ vectors of their respective values evaluated at the cell centers $\{\mathbf{x}_{1},...,\mathbf{x}_{N}\}$. For the inverse problem, we assume there are $N^{\text{obs}}_{y}$ observations of $y$, $\mathbf{y}^{\text{obs}}=\{y_{i}^{\text{obs}}\}_{i=1}^{N_{y}^{\text{obs}}}$, and $N_{u}^{\text{obs}}$ observations of $u$, $\mathbf{u}^{\text{obs}}=\{u^{\text{obs}}_{i}\}_{i=1}^{N_{u}^{\text{obs}}}$.
@@ 3.2 Hanford Site Case Study
The rPICKLE method is tested for estimating subsurface parameters at the Hanford Site, a United States Department of Energy site situated on the Columbia River in Washington State. We use Eqs. (102)–(104) to describe the two-dimensional (depth-averaged) groundwater flow at the Hanford Site. The ground truth conductivity and transmissivity fields are taken from a previous calibration study reported in Cole et al. (2001). Following Yeung et al. (2022), we employ the unstructured quadrilateral mesh with $N=1475$ cells (see Figure 1) to compute PDE residuals. The rPICKLE estimates of the posterior distribution of log-transmissivity are compared with that of the HMC and, for the mode of the distribution, with PICKLE.
@@
To enable a comparison between HMC and rPICKLE, we generate a lower-dimensional (smoother) $y$ field $y_{\mathrm{LD}}$ using iterative local averaging of $y_{\mathrm{HD}}$. Such averaging reduces the variance and increases the correlation length of the field Tartakovsky et al. (2017) and, therefore, reduces the dimensionality of the inverse problem. At the $k$th iteration, for FV element $i$, we calculate the geometric mean of $y^{(k)}_{j}$ over the adjacent $j$ elements and assign this value to $y^{(k+1)}_{i}$. Here, $y_{\mathrm{HD}}=y^{(k=0)}$ and $y_{\mathrm{LD}}=y^{(k=30)}$. We find that the ten-term CKLE of $y_{\mathrm{LD}}$ retains 95% of the total variance of this field.
@@
Figure 2: (a) The reference log-transmissivity field $y_{\mathrm{HD}}$ and (b) the reference hydraulic head field $u_{\mathrm{HD}}$ corresponding to $y_{\mathrm{HD}}$ obtained by solving Eqs. (102)–(104) with the finite volume method. (c) The low-dimensional field $y_{\mathrm{LD}}$ obtained by iterative local averaging of $y_{\mathrm{HD}}$ and (d) the hydraulic head field $u_{\mathrm{LD}}$ corresponding to $y_{\mathrm{LD}}$.
@@ 3.3 Prior Mean and Covariance Models of $y$ and $u$
Next, we generate $N^{\text{MC}}$ number of realizations of the stochastic $y$ field by independently sampling $\{\xi_{i}\}_{i=1}^{N_{\xi}}$ from the normal distribution and solving Eqs. (102)–(104) for each realization of $y$ using the FV method described in Yeung et al. (2022). The ensemble of $u$ solutions is used to compute the (ensemble) mean and covariance of $u$ using Eqs. (112) and (113). Then, the mean and covariance of $u$ are conditioned on $u$ observations using the GPR equations (114). Finally, the CKLE of $u$ in Eq. (4) is constructed by performing the eigenvalue decomposition of the conditional covariance of $u$, i.e., by solving the eigenvalue problem similar to Eq. (111). In this work, we set $N^{\text{MC}}=5,000$.
@@ 4.1 Low-Dimensional Problem
We first present results for the low-dimensional case with the ground truth $y$ field given by $y_{\mathrm{LD}}$. Here, we assume that 10 $y_{\mathrm{LD}}$ and $u_{\mathrm{LD}}$ observations are available. The locations of observations are randomly selected from the Hanford Site well locations. We use the HMC-sampled posterior distribution of $\bm{\xi}$ (and $y$), and the PICKLE-estimated MAP to benchmark the rPICKLE method with and without Metropolization. As stated earlier, we set $\sigma^{2}_{\xi}=\sigma^{2}_{\eta}=1$. The value of $\sigma^{2}_{r}$ is chosen to maximize the LPP of the posterior. However, the joint problem of minimizing the rPICKLE loss over $\bm{\xi}$ and $\bm{\eta}$ and maximizing the LPP over $\sigma^{2}_{r}$ is computationally challenging. Instead, we compute the inverse solutions for several values of $\sigma^{2}_{r}$ in the range $[10^{-5},1]$ and select the value producing the largest LPP. We also study the effect of $\sigma^{2}_{r}$ on uncertainty in the inverse solution and the performance of HMC and rPICKLE methods.
@@
In Figure 3, we also show the coordinates of the joint posterior mode given by the PICKLE solution. We can see that the coordinates of the modes of the marginal and bivariate distributions obtained from HMC and rPICKLE are in good agreement with the coordinates of the joint distribution mode. It should be noted that unless the posterior is Gaussian, the coordinates of the modes of marginal distributions and the corresponding coordinates of the joint distribution mode may not coincide. Because the coordinates of the marginal and joint distribution modes are close to each other, this indicates that the posterior distribution is close to Gaussian.
@@
Finally, we present the Bayesian predictions of $y$. For each of the three methods and $\sigma^{2}_{r}=10^{-2}$, Figure 4 shows the estimated mean $\mu_{\hat{y}}(\mathbf{x}|\mathcal{D}_{res})$ field, the absolute difference between the mean and the reference field $|y-\mu_{\hat{y}}(\mathbf{x}|\mathcal{D}_{res})|$, the standard deviation $\sigma_{\hat{y}}(\mathbf{x}|\mathcal{D}_{res})$, and the coverage plot showing the locations where the reference solution is within the $95\%$ confidence interval. In Table 1, we summarize results in terms of the relative $\ell_{2}$ error and $\ell_{\infty}$ error between the predicted mean and the reference $y_{\mathrm{LD}}$ fields, LPP, and the coverage (percentage of nodes where the reference solution is within the $95\%$ confidence interval) for $\sigma^{2}_{r}=10^{-5}$, $10^{-4}$, $10^{-2}$, $10^{-1}$, and $10^{0}$. For a given value of $\sigma^{2}_{r}$, all three methods produce a posterior mean close to the PICKLE’s estimate of MAP, and similar LPPs and coverages. This indicates that (i) the HMC and rPICKLE sampled distributions converge to the same posterior (note that in Section 2.4.1, we only prove the consistency of rPICKLE in the linear case), and (ii) Metropolization makes the posterior more descriptive of the reference field, i.e., it yields larger LPP and smaller $\ell_{2}$ error, but the improvements are less than $1\%$. We also find that, for this problem, the total acceptance rate for Metropolization is above $95\%$, which agrees with the findings in Oliver et al. (1996); Wang et al. (2018) that the acceptance rate in randomized algorithms is, in general, very high. It should be noted that the Metropolization requires the estimation of the Jacobian, which becomes computationally expensive for high-dimensional problems. For this reason, in the high-dimensional case presented in Section 4.2, we do not perform Metropolization and accept all samples generated by the rPICKLE algorithm.
@@
According to Table 1, for a given value of $\sigma^{2}_{r}$, HMC and rPICKLE yield similar mean estimates of $y$ in terms of the point errors and the relative $\ell_{2}$ errors. Both, the HMC and rPICKLE errors are smallest for $\sigma^{2}_{r}=10^{-2}$. We find that the LPPs in these methods are also the largest for this value of $\sigma^{2}_{r}$. This indicates that $\sigma^{2}_{r}=10^{-2}$ provides the most informative posterior distribution of $y$. It is worth reminding that $\sigma^{2}_{r}=\gamma$, the regularization parameter in PICKLE. Because PICKLE’s $\ell_{2}$ and $\ell_{\infty}$ errors are also smallest for $\sigma^{2}_{r}=10^{-2}$, we conclude that this value of $\sigma^{2}_{r}$ also provides the optimal regularization for this problem.
@@
Figure 4: The low-dimensional $y$ field estimated from (a) HMC, (b) rPICKLE, and (c) rPICKLE with Metropolization. The first column shows the posterior mean estimates; the second column shows the point errors in the predicted mean with respect to the reference $y_{LD}$ field; the third column shows the posterior standard deviation of $y$; and the fourth column shows the coverage of the reference field with the $95\%$ credibility interval, where the zero and one values correspond to the reference field being outside and inside the credibility interval, respectively.
Table 1: The summary of the rPICKLE, HMC, and PICKLE results for the low-dimensional problem with different values of $\sigma^{2}_{r}$. Shown are the relative $\ell_{2}$ and $\ell_{\infty}$ errors in the estimated $y$ field with respect to the reference field, the LPP, and the percentage of coverage of $y_{\mathrm{LD}}$ by the $95\%$ confidence interval. Note that the PICKLE solution only provides MAP, and LPP and coverage for the PICKLE solution are not given here.
@@ 4.2 High-Dimensional Problem
Here, we consider the IUQ problem where the $y$ field is given by $y_{\mathrm{HD}}$. First, we estimate the posterior distributions for $\sigma^{2}_{r}=10^{-4}$, $10^{-2}$, and $10^{-1}$ given $N_{y}^{\text{obs}}=100$ observations of the $y_{\mathrm{HD}}$ field. Later, we estimate posteriors for $N_{y}^{\text{obs}}=50$ and $200$ to study the dependence of the posterior on the number of $y$ measurements. In all simulations in this section, we assume that $N_{u}^{\text{obs}}=323$, i.e., $u$ measurements are available at all wells.
@@
Table 2: The summary of the rPICKLE and PICKLE results for the high-dimensional problem for the priors based on different numbers of $y$ observations and $\sigma^{2}_{r}$. Shown are the relative $\ell_{2}$ and $\ell_{\infty}$ errors in the estimated $y$ field with respect to the reference field, the LPP, and the percentage of coverage of $y_{\mathrm{LD}}$ by the $95\%$ confidence interval.
@@
Table 2 summarizes the relative $\ell_{2}$ and $\ell_{\infty}$ errors, LPP, and coverage in rPICKLE estimates of $y$ for $\sigma^{2}_{r}=10^{-4}$, $10^{-2}$, and $10^{-1}$. We find that the smallest rPICKLE errors and the largest LPP are achieved for $\sigma^{2}_{r}=10^{-2}$. However, we find that LPP is more sensitive to $\sigma^{2}_{r}$ than to $\ell_{2}$ error–$\ell_{2}$ errors vary by less than 7% for the considered $\sigma^{2}_{r}$ values, while LPP values change by more than 100%. Also, the $\ell_{2}$ error is 1% smaller for $\sigma^{2}_{r}=10^{-4}$ than for $\sigma^{2}_{r}=10^{-1}$ but the LPP is 70% larger for $\sigma^{2}_{r}=10^{-1}$ than for $\sigma^{2}_{r}=10^{-4}$. Therefore, we conclude that LPP provides a better criterion for selecting $\sigma^{2}_{r}$ than $\ell_{2}$ errors. Errors in the PICKLE and rPIKCLE predictions of $y$ are very similar, and $\sigma^{2}_{r}=10^{-2}$ also provides the best value of the regularization coefficient for PICKLE, i.e., the PICKLE error is smallest for $\gamma=\sigma^{2}_{r}=10^{-2}$.
@@
Figure 6 shows the rPICKLE estimate of the posterior mean of $y$, the absolute point difference between the mean and reference $y_{\mathrm{HD}}(x)$ fields, the posterior standard deviation of $y(x)$, and the coverage for $\sigma^{2}_{r}=10^{-2}$ and $\sigma^{2}_{r}=10^{-4}$. We see significant differences in rPICKLE predictions for different $\sigma^{2}_{r}$. Errors in the predictions with $\sigma^{2}_{r}=10^{-4}$ are in general larger than in the predictions with $\sigma^{2}_{r}=10^{-2}$ with the maximum point error being 50% larger. As expected, the posterior standard deviations are generally larger in the prediction with the larger $\sigma^{2}_{r}$. However, the maximum point standard deviation is larger in the simulation with the smaller $\sigma^{2}_{r}$. We also see that $\sigma^{2}_{r}=10^{-2}$ produces a better coverage–the reference $y_{\mathrm{HD}}$ field is within the confidence interval in 82% of all predicted locations versus 65% for $\sigma^{2}_{r}=10^{-2}$. We reiterate that the LPP for $\sigma^{2}_{r}=10^{-4}$ is -2792, which is significantly smaller than that for $\sigma^{2}_{r}=10^{-2}$ (2032).
Next, we study uncertainty in the inverse rPICKLE solution as a function of $N_{y}^{\text{obs}}$. Figure 7 depicts the estimates of $y$ obtained with rPICKLE for $N_{y}^{\text{obs}}=50$ and $200$ and $\sigma^{2}_{r}=10^{-4}$. The estimates for $N_{y}^{\text{obs}}=100$ are given in Figure 6. Table 2 summarizes the relative $\ell_{2}$ and $\ell_{\infty}$ errors, LPP, and the percent of coverage of the corresponding posteriors. As $N_{y}^{\text{obs}}$ increases, the posterior mean becomes closer to $y_{\mathrm{HD}}$, and the posterior variance of $y$ decreases. The LPP increases with $N_{y}^{\text{obs}}$, indicating that the posterior distribution becomes more informative. Also, we see that for all values of $N_{y}^{\text{obs}}$, the coverage for the $y_{\mathrm{HD}}$ field is good, with the best coverage ($75\%$) achieved for $N_{y}^{\text{obs}}=200$.